OpenStudy (nick88888888):

What is the range of y = 5x -2 if the domain is {-3, -1, 0, 1, 3} OuO

OpenStudy (haseeb96):

just put domain number one by one in the given function and get the range for example here i put -3 in y=5x-2 in the place of x y=5(-3)-2 y=-15-2 y=-17 is the range of function here y is the function and now the remaining range is yur job now find them

OpenStudy (nick88888888):

@Haseeb96 Oh Thank you ^_^, That was a big help T^T

OpenStudy (haseeb96):

welcome by brother ^_^